OpenStudy (luigi0210):

Use cylindrical shells to find the volume:

OpenStudy (luigi0210):

xy=1, x=0, y=1, y=3

ganeshie8 (ganeshie8):

whats the axis of revolution ?

OpenStudy (anonymous):

x - axis or y - axis?

OpenStudy (luigi0210):

About the x-axis, sorry about that

OpenStudy (anonymous):

\[\int\limits_{a}^{b} 2 \pi (shell radius)(shell height) dy\]

OpenStudy (luigi0210):

Teach me batman :3

OpenStudy (anonymous):

So it's from 1 to 3  \[\int\limits_{1}^{3} 2 \pi (y) \frac{ 1 }{ y }dy \] \[= \int\limits_{1}^{3} 2\pi (1) dy = 2\pi(y) 1 ~ \to ~ 3 = 4\pi\]

OpenStudy (anonymous):

I gtg now, good luck lui :p

OpenStudy (luigi0210):

Thank you batman :D